# Optimization of modulation transfer protocol for Rydberg RF receivers

### v1

@@ IV Expansion of RF bandwidth
In order to examine the performance of the MTP in comparison to the CP, we focus on weak and detuned RF fields. In particular, we try to see whether the MTP permits to expand the RF bandwidth. The sensitivity $S$ [schlossbergerRydbergStatesAlkali2024], expressed in $\mathrm{V.m^{-1}.{Hz}^{-1/2}}$ is defined as
@@
Finally, we wish to determine in which range of parameters the modulation transfer protocol outperforms the conventional protocol and to what extent. To perform this comparison, since the noise associated with the two protocols is similar, we calculate the ratio between the slopes of the system responses in the two protocols. This ratio is shown in color maps in Fig.9, both experimentally and theoretically. The white lines indicate where the sensitivity levels for both protocols are equal, i.e., when the sensitivity improvement factor equals 1. Above that line, the modulation transfer protocol outperforms the conventional one. In the experiment (see Fig.9(a)), the modulation transfer protocol exhibits better performance as soon as $\Delta_{\mathrm{RF}}/2\pi=3$ MHz for vanishing RF field amplitudes. Starting from $\Delta_{\mathrm{RF}}/2\pi=4$ MHz, the improvement in sensitivity exceeds 2. This limit closely follows the RF bandwidth at -6 dB. For even greater RF detunings, this improvement goes beyond a factor of 20. As discussed previously, the simulations of Fig. 9(b) closely replicate the experiment.
@@
Table 2: Sensitivities of the RF field amplitude measurement for the two measurement protocols and for several values of the RF detuning and $E_{\mathrm{RF}}=0.05\,\mathrm{V/m}$. The sensitivities are given in $\mathrm{\mu V.cm^{-1}.Hz^{-1/2}}$. The first (second) part of the table is based on a noise value $V_{0}$ measured experimentally (deduced from the shot noise limit).
@@
We finally give in this last paragraph a comparison between the sensitivities, defined by Eq. (11), of the two protocols. Table 2 presents the sensitivities of the two protocols for various values of the RF detuning. For each case, we give two values of the sensitivity: one is obtained by taking the shot noise limit for $V_{0}$ in Eq. (11), while the other is based on measurements of the noise $V_{0}$ in our experiment. In both cases, the slope $\partial V/\partial E_{\mathrm{RF}}$ is measured. We notice that the MTP outperforms the CP as soon as $\Delta_{\mathrm{RF}}/2\pi>3\,\mathrm{MHz}$ for weak fields. Moreover, the improvement of the sensitivity for the MTP with respect to the CP is even better when we consider the measured noise than when we consider the shot noise limit. This is due to the fact that the MTP corresponds to a measurement at 3 MHz, where the detected noise is much lower than at the lower frequency (100 kHz) where the CP is implemented. Moreover, we observe that these sensitivity measurements compare very favorably with those in the existing literature [fanAtomBasedRF2015, kumarAtomBasedSensingWeak2017] within the same range of RF frequencies.
@@ V Conclusion
The two protocols are thus complementary: the conventional protocol is recommended for RF field detunings smaller than 3 MHz, while the modulation protocol should be used for larger detunings. Since both protocols use the same apparatus, switching between them is straightforward and can be accomplished by modifying the modulation applied to the coupling beam AOM. One can also operate both protocols simultaneously within the same vapor cell. The optimized modulation parameters presented in this paper only apply to Rubidium 85; they should be tailored to any other Alkali atom and adapted when the number of probed atoms is modified. However, it is important to note that this technique does not achieve RF bandwidths or sensitivities comparable to alternative techniques based on the use of an RF local oscillator [andersonContinuousfrequencyMeasurementsHighintensity2017, huContinuouslyTunableRadio2022, yaoSensitivityEnhancementFardetuned2022, liRydbergAtombasedAM2022, simonsContinuousRadiofrequencyElectricfield2021]. Nonetheless, the protocol presented in this article is an all-optical method that does not require any external RF antenna, which would undermine the goal of developing a dielectric sensor.
@@ Acknowledgements.
This work was achieved in spite of the lack of support from ENS Paris-Saclay. We acknowledge the significant contributions of Joseph Delpy and Nadia Belabas, whose insightful discussions have profoundly impacted our work. Additionally, we recognize Sébastien Rousselot for his vital technical support. This work was funded by the French Defense Innovation Agency, Quantum Saclay, Agence Nationale de la Recherche (ANR) for France 2030 under Grant Agreement No. ANR-23-PETQ-0004 and the European Defense Fund (EDF) under Grant Agreement No. EDF-2021-DIS-RDIS-ADEQUADE (No. 101103417). This study was funded by the European Union. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@ IV Expansion of RF bandwidth
This section explores the expansion of the RF bandwidth allowed by the MTP optimization performed in Section III.B. We will first examine the bandwidth expansion around a single transition. Next, we will compare the sensitivities of the two protocols. Finally, we will assess whether the bandwidth improvement is sufficient to bridge the frequency gap between two Rydberg transitions.
IV.1 Expansion of RF Bandwidth around a Single Transition
In order to examine the performance of the MTP in comparison to the CP, we focus on weak and detuned RF fields. In particular, we try to see whether the MTP permits to expand the RF bandwidth. The sensitivity $S$ [29], expressed in $\mathrm{V.m^{-1}.{Hz}^{-1/2}}$ is defined as
@@ IV.1 Expansion of RF Bandwidth around a Single Transition
IV.2 Sensitivity Comparison between the Two Protocols
Finally, we wish to determine in which range of parameters and to which extent the modulation transfer protocol outperforms the conventional protocol.
To perform this comparison, since the noise associated with the two protocols is similar, we calculate the ratio between the slopes of the system responses in the two protocols. This ratio is shown in color maps in Fig.9, both experimentally and theoretically. The white lines indicate where the sensitivity levels for both protocols are equal, i.e., when the sensitivity improvement factor equals 1. Above that line, the modulation transfer protocol outperforms the conventional one. In the experiment (see Fig.9(a)), the modulation transfer protocol exhibits better performance as soon as $\Delta_{\mathrm{RF}}/2\pi=3$ MHz for vanishing RF field amplitudes. Starting from $\Delta_{\mathrm{RF}}/2\pi=4$ MHz, the improvement in sensitivity exceeds 2. This limit closely follows the RF bandwidth at -6 dB. For even greater RF detunings, this improvement goes beyond a factor of 20. As discussed previously, the simulations of Fig. 9(b) closely replicate the experiment.
@@ IV.2 Sensitivity Comparison between the Two Protocols
[…]
@@
We finally give in this last paragraph a comparison between the sensitivities, defined by Eq. (11), of the two protocols. Table 2 presents the sensitivities of the two protocols for various values of the RF detuning. For each case, we give two values of the sensitivity: one is obtained by taking the shot noise limit for $V_{0}$ in Eq. (11), while the other is based on measurements of the noise $V_{0}$ in our experiment. In both cases, the slope $\partial V/\partial E_{\mathrm{RF}}$ is measured. We notice that the MTP outperforms the CP as soon as $\Delta_{\mathrm{RF}}/2\pi>3\,\mathrm{MHz}$ for weak fields. Moreover, the improvement of the sensitivity for the MTP with respect to the CP is even better when we consider the measured noise than when we consider the shot noise limit. This is due to the fact that the MTP corresponds to a measurement at 3 MHz, where the detected noise is much lower than at the lower frequency (100 kHz) where the CP is implemented. Moreover, we observe that these sensitivity measurements compare very favorably with those in the existing literature [9, 17] within the same range of RF frequencies.
IV.3 Closing a Frequency Gap between Two Rydberg Transitions
Now that we know that the MTP expands the RF bandwidth around a single transition, let us examine whether it can bridge the gap between two nearby Rydberg transitions.
[…]
@@ V Conclusion
The two protocols are thus complementary: the conventional protocol is recommended for RF field detunings smaller than 3 MHz, while the modulation protocol should be used for larger detunings. Since both protocols use the same apparatus, switching between them is straightforward and can be accomplished by modifying the modulation applied to the coupling beam AOM. One can also operate both protocols simultaneously within the same vapor cell. The optimized modulation parameters presented in this paper only apply to Rubidium 85; they should be tailored to any other Alkali atom and adapted when the number of probed atoms is modified. We demonstrated that the modulation transfer protocol enables closing RF frequency gaps of at least 166 MHz, which is promising for developing a frequency-continuous sensor. The results presented in this article could be improved by heating the cell, using optical repumping, or increasing the atoms’ transit time. However, it is important to note that this technique does not achieve RF bandwidths or sensitivities comparable to alternative techniques based on the use of an RF local oscillator [2, 14, 40, 19, 33]. It is also not sensitive to the phase of the RF field. Nonetheless, the protocol presented in this article is an all-optical method that does not require any external RF antenna, which would undermine the goal of developing a dielectric sensor.
@@ Acknowledgements.
We acknowledge the significant contributions of Joseph Delpy and Nadia Belabas, whose insightful discussions have profoundly impacted our work. Additionally, we recognize Sébastien Rousselot for his vital technical support. This work was funded by the French Defense Innovation Agency, Quantum Saclay, Agence Nationale de la Recherche (ANR) for France 2030 under Grant Agreement No. ANR-23-PETQ-0004 and the European Defense Fund (EDF) under Grant Agreement No. EDF-2021-DIS-RDIS-ADEQUADE (No. 101103417). This study was funded by the European Union. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them.